OpenStudy (abmon98):

Integrate 2x-1/36+(x-3)^2

OpenStudy (abmon98):

2x-1/36(1+(x-3)^2/36) u=x-3/6 u^2=(x-3)^2/36 12u+5=(2x-6)+5 12u+5/36(1+u^2)

OpenStudy (anonymous):

break in to two parts

OpenStudy (anonymous):

one will give an arctangent, the other a log

OpenStudy (anonymous):

assuming it is \[\int \frac{2x-1}{36+(x-3)^2}dx\]

OpenStudy (abmon98):

\[\int\limits_{}^{}12u/36(1+u^2)+\int\limits_{}^{}5/36(1+u^2)\]

OpenStudy (abmon98):

and yes its in this form @satellite73

OpenStudy (mathmale):

Which integral will result in an inverse tangent function? Can you perform this integration? Which method would you use to integrate the other part of this integral problem?

OpenStudy (abmon98):

=1/6ln(1+u^2)+5/36tan^-1(u)

OpenStudy (mathmale):

I've checked and found that your first integral is fine.

OpenStudy (mathmale):

Next time you want verification of correctness, please show all of your work as part of the deal.

OpenStudy (zarkon):

what was your substitution for \(dx\)

OpenStudy (abmon98):

Okay I am sorry, but my answer is incorrect according to the answer sheet.

OpenStudy (abmon98):

So here is what i did i factored out the 36 from the denominator and used the susbtitution method where u=x-3/6 therefore u^2=(x-3)^2/36. I then rewrote the expression u=x-3/6 to be 12u+5 to cope with the numerator. 12u+5/36(1+u^2) did a split to the fraction and then integrated both 12u/36(1+u^2)+5/36(1+u^2)

OpenStudy (zarkon):

if \[u=\frac{x-3}{6}\] then \[6u=x-3\] and \[6du=dx\]

OpenStudy (abmon98):

Oh thanks I just realised my mistake here i forgot the du part

OpenStudy (zarkon):

you also need to replace u in the end. your final answer should be a function of x

OpenStudy (abmon98):

Here is my final answer after i have altered some stuff ln(1+(x-3/6)^2)+5/6tan^-1(x-3/6)

OpenStudy (zarkon):

+c